Ex 6.3 Class 6 Maths Question 4.
Find :
(a) (-7) – 8 – (-25)
(b) (-13) + 32 – 8 – 1
(c) (-7) + (-8) + (-90)
(d) 50 – (-40) – (-2)

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) (-7) – 8 – (-25) = (-7) – 8 + 25 [ ∵ Additive inverse of – 25 is 25] = -7 + 17 = -7 + 7 +10 [∵ (-a) + (+a) = 0] = 0 + 10 = 10. (b) (-13) + 32 – 8 – 1 = (-13) + (13) + 19 – (8 + 1) = 0 + 19 – 9 = 19 – 9 [∵ (-13) + (13) = 0] = 10 + 9 – 9 = 10 + 0 = 10. [(+9) – (+9) = 0] (c) (-7) + (-8) + (-90) = – (7 + 8) + (-90) = -15 + (-90) = -(15 + 90) = -105. (d) 50 – (-40) – (-2) = 50 – [- 40 – 2] = 50 – (-42) = 50 + 42 = 92.